OpenStudy (superhelp101):

How many moles of sulfuric acid are needed to produce 57.8 milliliters of water? Water has the density of 0.987 g/mL. Show all steps of your calculation as well as the final answer. NaOH + H2SO4 → Na2SO4 + H2O

OpenStudy (superhelp101):

@Abhisar

OpenStudy (superhelp101):

do i have to balance the equation first?

OpenStudy (abhisar):

1 mL of water = 0.987 gram of water

OpenStudy (abhisar):

yes balancing is must

OpenStudy (abhisar):

now calculate the amount of water in grams for 57.8 milliliters

OpenStudy (superhelp101):

balancedd equation: H2SO4 + 2 NaOH -> Na2SO4 + 2 H2O grams: 57.0486

OpenStudy (superhelp101):

ok

OpenStudy (abhisar):

according to the balanced equation 1 mole of H2SO4 produces 2 moles of water (18 \(\times\)2=36grams). ok ?

OpenStudy (superhelp101):

how do i solve that?

OpenStudy (abhisar):

solve what ?

OpenStudy (superhelp101):

i mean what do i do after i found in grams and the balaced equation

OpenStudy (abhisar):

i am explaining u that only. Respond to my questions. U got the above line ?

OpenStudy (superhelp101):

yea

OpenStudy (abhisar):

Now density of water is given 0.987 g/mL, so 1 mL of water = 0.987 gram. Thus, 57.8 mL = 0.987\(\times\)57.8=57.086 gram

OpenStudy (superhelp101):

alright

OpenStudy (abhisar):

Now since 1 mole of H2SO4 produces 36 gram of water how many moles will produce 57.086 gram water ?

OpenStudy (abhisar):

Find it, that will be the answer

OpenStudy (superhelp101):

1.58?

OpenStudy (abhisar):

\(\huge\checkmark\)

OpenStudy (superhelp101):

Thanks! u da bomb <3

OpenStudy (abhisar):

Scared!!

OpenStudy (superhelp101):

huh?

OpenStudy (abhisar):

BOMB!

OpenStudy (abhisar):

Just kiddin !

OpenStudy (superhelp101):

WOW LOL